tax:

An Algebra 2 question that I want to see if I got correct:

tax:

\[5x^{2/3} + 14 = 59\] And my final answer was: \[x = 27\] Is this correct? @xxaikoxx

tax:

Because I know that there was a point in the problem where you had to find the reciprocal of the fraction as well, so my answer that I got depends on how I did those steps

tax:

@xxaikoxx Here

xXAikoXx:

@tax wrote:
\[5x^{2/3} + 14 = 59\] And my final answer was: \[x = 27\] Is this correct? @xxaikoxx
Unfortunately, your answer is not correct. Here's why: Subtract from both sides here \[\frac{ 5x^2}{3} = 45\] Then multiply both sides by 3 \[5x^2 = 135\] Next, divide by 5 \[x^2 = 27\] 4. Now we take the square root \[x=±\sqrt{27} =±3\sqrt{3} \] Therefore, x = 27 is NOT true because it is the answer to \[x^2\] Instead, ** x = 3. ** This will right.

tax:

@xxaikoxx wrote:
@tax wrote:
\[5x^{2/3} + 14 = 59\] And my final answer was: \[x = 27\] Is this correct? @xxaikoxx
Unfortunately, your answer is not correct. Here's why: Subtract from both sides here \[\frac{ 5x^2}{3} = 45\] Then multiply both sides by 3 \[5x^2 = 135\] Next, divide by 5 \[x^2 = 27\] 4. Now we take the square root \[x=±\sqrt{27} =±3\sqrt{3} \] Therefore, x = 27 is NOT true because it is the answer to \[x^2\] Instead, ** x = 3. ** This will right.
Yay, thank you!

tax:

I have like 4 more by the way :sob: